OpenStudy (anonymous):

Solve P = 2(l + w) for l. What are the missing values in the table? (1 point) P w l 14 2 ? 22 8 ? Create your own unique literal equation and solve for one of the variables. Show your work. Then, using complete sentences, explain how you solved for the variable you chose. (2 points) Using complete sentences, explain how solving a literal equation is similar to or different from simplifying an expression such as 6 - 2(52 + 7) � 4. (2 points) Using complete sentences, explain what might happen if the order of operations was used to solve a literal equation. (1 point)

OpenStudy (anonymous):

I=(P-2W)/2

OpenStudy (anonymous):

plug in the values and solve for I

OpenStudy (anonymous):

the perimeter is 72.5' width is 16.5' I need the length of the wall

OpenStudy (anonymous):

Well the equation given to find the perimiter "P" is P=2(L+W). If we rearange this equation to solve for L we get L=(P-2W)/2. Then you just have to plug in your Perimeter and Width values.

OpenStudy (anonymous):

The first one would look like this in your calculator: (14-(2x2))/2

OpenStudy (anonymous):

i need to show 2 ways to find the length.

OpenStudy (anonymous):

No the question is asking for you to solve the equation given for the variables. In this case the variables are P,L, and W. Being that you are solving for the length from the perimeter there is only one way to solve for L. The second part of the question asks you to explain how solving for variables in an equation is the same as solving for numbers

OpenStudy (anonymous):

Nooooo. Henry is putting a new baseboard around his room. He used the formula p=2(l+w) to find the perimeter. The perimeter is 72.5'. He remembers the width is 16.5'. Show 2 different ways to find the length of the other wall.